OpenStudy (anonymous):

Simplify completely 18 s t to the fourth power over 52 s cubed t times 16 s cubed over 9 s squared t A 8 t squared over 13 s B 8 t over 13 s squared C 4 t squared over 13 s D 4 t over 39 s squared

OpenStudy (jdoe0001):

\(\bf \cfrac{18st^4}{52s^3t} \times \cfrac{16s^3}{9s^2} \implies \cfrac{18stttt}{52ssst} \times \cfrac{16sss}{9sss}\) see any like-terms to cancel out there?

OpenStudy (anonymous):

16SSS 9SSS?

OpenStudy (jdoe0001):

\(\bf s\times s\times s = s^3\\ t \times t\times t\times t = t^4\) and so on

OpenStudy (jdoe0001):

hmm, I put an extra "s" in the 9, lemme fix that

OpenStudy (jdoe0001):

\(\bf \cfrac{18st^4}{52s^3t} \times \cfrac{16s^3}{9s^2} \implies \cfrac{18stttt}{52ssst} \times \cfrac{16sss}{9ss}\\ \implies \cfrac{(9)(2)stttt}{(4)(13)ssst} \times \cfrac{(4)(4)sss}{9ss}\)

OpenStudy (anonymous):

LOST

OpenStudy (jdoe0001):

well.... ok, what is the answer for say \(\bf \cfrac{2}{2}\)

OpenStudy (anonymous):

1

OpenStudy (jdoe0001):

what about \(\bf\cfrac{4}{2}\) ?

OpenStudy (anonymous):

YES